Solve
x^4-8x^3+21x^2-20x+5=0
given that the sum of two of its roots is 4 .
1
Expert's answer
2018-06-27T04:06:08-0400
x^4-8x^3+21x^2-20x+5=0

(x^2-5x+5)(x^2-3x+1)=0

x^2 -5x + 5 = 0

x = 3/2 +- sqrt(5)/2

x^2-3x+1 = 0

x = 5/2 +- sqrt(5)/2